OpenStudy (anonymous):

what is the radius of a circelwith an area of 350ft^2

OpenStudy (anonymous):

area of a circle is \[\Pi r^2\] so equate \[\Pi r^2=350ft^2\] divide by pi and take the square root

OpenStudy (anonymous):

well on my paper it says the equation is sqrt A/pi= r

OpenStudy (anonymous):

thats the same thing.\[\sqrt{A/\Pi}=r\] is the same as dividing the area "A" by Pi and then taking the square root. so A=350, divide by Pi, and then take the square root of that answer

OpenStudy (anonymous):

so about 10.55?

OpenStudy (anonymous):

yep thats it!

OpenStudy (anonymous):

thanks. :] can you help me some more?

OpenStudy (anonymous):

i might? what the question?

OpenStudy (anonymous):

let f(x) = 2x +5 find the value of (f 0 f ^-1) (-1/2)

OpenStudy (anonymous):

i'm not sure if i understand the last part, can you try rewriting it?

OpenStudy (anonymous):

is it a function composition?

OpenStudy (anonymous):

well i'm going on to a different problem, post your question on the main part

OpenStudy (anonymous):

well its like f of f^-1

OpenStudy (anonymous):

than that's what it is. first find the inverse of the function 2x+5. then plug in -1/2 into the inverse function. then we will take f(x)=2x+5 and replace the x with the answer we got for inverse.

OpenStudy (anonymous):

so for the inverse we replace y with x. so it'll be changed to x=2y+5. solve for y. y=(-5x+5)/2. plug in (-1/2) and we get y=-5(1/2)/2. plug in this answer into y=2x+5 and we get y=2(-5(1/2))/2+5. the 2's cancel and we get -5(1/2)+5. this equals (-1/2).

OpenStudy (anonymous):

thank you.! so much

OpenStudy (anonymous):

no problem!